Determine projau the orthogonal projection of u and a and deduce ||projau|| for (2.1) u =<−1,3 >, a=<−1,−3>;

(2.2) u =<−2,1,−3 >, a =<−2,1,2 >.


1
Expert's answer
2021-06-25T07:15:01-0400

(2.1) "\mathbf{u} =\langle -1 ,\ 3\rangle ,\quad \mathbf{a} =\langle-1,\ -3\rangle"


"\text{proj}_ \mathbf{a} \mathbf{u} =\big(\frac{ \mathbf{u}\ \cdot \ \mathbf{a}}{\| \mathbf{a}\|^2} \big)\ \mathbf{a}" and "\| \text{proj}_ \mathbf{a} \mathbf{u}\|=\frac{|\mathbf{u}\ \cdot \ \mathbf{a}|}{\| \mathbf{a}\|}"


"\mathbf{u}\ \cdot \ \mathbf{a}=1-9=-8"

"\| \mathbf{a}\|^2=1+9=10"


"\text{proj}_ \mathbf{a} \mathbf{u} =-\frac{8}{10} \mathbf{a}=\langle0.8,\ 2.4\rangle" and "\|\text{proj}_ \mathbf{a} \mathbf{u}\|=\frac{8}{\sqrt{10}}=\sqrt{6.4}"


(2.2) "\mathbf{u} =\langle -2,1 ,\ - 3\rangle ,\quad \mathbf{a} =\langle-2,1,\ 2\rangle"


"\text{proj}_ \mathbf{a} \mathbf{u} =\big(\frac{ \mathbf{u}\ \cdot \ \mathbf{a}}{\| \mathbf{a}\|^2} \big)\ \mathbf{a}" and "\| \text{proj}_ \mathbf{a} \mathbf{u}\|=\frac{ |\mathbf{u}\ \cdot \ \mathbf{a}|}{\| \mathbf{a}\|}"


"\mathbf{u}\ \cdot \ \mathbf{a}=4+1-6=-1"

"\| \mathbf{a}\|^2=4+1+4=9"


"\text{proj}_ \mathbf{a} \mathbf{u} =-\frac{1}{9} \mathbf{a}=\langle
\frac{2}{9},\ -\frac{1}{9},\ -\frac{2}{9}
\rangle" and "\|\text{proj}_ \mathbf{a} \mathbf{u}\|=\frac{1}{\sqrt{9}}=\frac{1}{3}"


Answer:

(2.1) "\text{proj}_ \mathbf{a} \mathbf{u}=\langle0.8,\ 2.4\rangle" and "\|\text{proj}_ \mathbf{a} \mathbf{u}\|=\sqrt{6.4}" 

(2.2) "\text{proj}_ \mathbf{a} \mathbf{u} =\langle
\frac{2}{9},\ -\frac{1}{9},\ -\frac{2}{9}
\rangle" and "\|\text{proj}_ \mathbf{a} \mathbf{u}\|=\frac{1}{3}"